# Hom-orthogonal modules and brick-Brauer-Thrall conjectures

### v1

@@ 4. Orthogonal orbits and irreducible components
In the following, $A$ is always assumed to be a basic, connected finite dimensional algebra of rank $n$ over an algebraically closed field $k$. For an irreducible component $\mathcal{Z}\in\operatorname*{Irr}(A)$, recall that two orbits $O_{X}$ and $O_{Y}$ in $\mathcal{Z}$ are said to be orthogonal if $X$ and $Y$ are Hom-orthogonal modules. In this section we focus on the behaviour of the Hom-orthogonal modules in the irreducible components of $A$. Through this geometric approach, we obtain a surprisingly elementary equivalent condition for the existence of an infinite family of bricks of the same dimension (Theorem 4.2 and Corollary 4.3). In particular, we prove Theorem B from Section 1. We also give a characterization of brick-finite algebras as a generalization of the local algebras.
We begin with a lemma which is handy in the rest.
@@
Note that $\mathcal{Z}$ has a dense orbit if and only if $\mathcal{Z}=\overline{O_{M}}$, for some $M\in\mathcal{Z}$, that is, $O_{M}$ is an open orbit in $\mathcal{Z}$. Let $\mathcal{Z}$ contain a pair of modules $X$ and $Y$ with $\operatorname{Hom}_{A}(X,Y)=0$. Observe that the function ${\rm dim}_{k}\operatorname{Hom}_{A}(X,-)$ is upper-semicontinuous and attains the minimal value of 00 in $\mathcal{Z}$, particularly because we assumed $\operatorname{Hom}_{A}(X,Y)=0$. Therefore, there exists a non-empty open set $O$ in $\mathcal{Z}$ where that minimum is attained. But the open set $O$ intersects $O_{M}$, yielding $\operatorname{Hom}_{A}(X,M)=0$. On the other hand, we know that $O_{X}\subseteq\mathcal{Z}=\overline{O_{M}}$, meaning that $M$ degenerates to $X$. Thus, by Theorem 2.2, we have a short exact sequence $0\rightarrow X\rightarrow M\oplus Z\rightarrow Z\rightarrow 0$ in $\operatorname*{mod}A$, for some $Z$. Apply $\operatorname{Hom}_{A}(-,M)$ to this short exact sequence and note that, from $\operatorname{Hom}_{A}(X,M)=0$, we obtain an isomorphism $\operatorname{Hom}_{A}(Z,M)\simeq\operatorname{Hom}_{A}(Z,M)\oplus\operatorname{Hom}_{A}(M,M)$. Consequently, $\operatorname{Hom}_{A}(M,M)=0$, which is clearly a contradiction. ∎
The preceding lemma leads to the following result. In particular, we verify the Strong Semibrick Conjecture for the algebras under consideration (see Remark 3.5).
Theorem 4.2
@@
The previous theorem implies two interesting corollaries that we mention below. The first one gives a new characterization of the brick-finite algebras as a generalization of the local algebras from the viewpoint of geometry. Since the result follows from the above theorem, we eliminate the proof. However, to motivate the assertion, let us recall that if $A$ is a local algebra, then for every pair of nonzero $A$-modules $X$ and $Y$, we have $\operatorname{Hom}_{A}(X,Y)\neq 0$.
@@
We now state another corollary of Theorem 4.2. In light of Proposition 3.4, this gives a new equivalent condition for the $2nd$ bBT conjecture. To put this result in perspective, recall that the celebrated Second Brauer-Thrall conjecture (now theorem) asserts that $A$ is a representation-infinite algebra if and only if there are infinitely many $d\in\mathbb{Z}_{>0}$ for which $\operatorname*{ind}(A,d)$ is an infinite set. The following result implies that, to verify the $2nd$ bBT conjecture for $A$, it suffices to show that for some $d\in\mathbb{Z}_{>0}$ and infinitely many pairs $X$ and $Y$ in $\operatorname*{ind}(A,d)$, we have $\operatorname{Hom}_{A}(X,Y)=0$.
Corollary 4.4
For an algebra $A$, the following are equivalent:
  1. (1)
    $A$ admits an infinite family of bricks of the same dimension;
  2. (2)
    There is an infinite family $\{X_{i}\}_{i\in I}$ of $A$-modules of dimension $d$ such that, for distinct pairs $i,j\in I$, either $\operatorname{Hom}_{A}(X_{i},X_{j})=0$ or $\operatorname{Hom}_{A}(X_{j},X_{i})=0$.
Proof.
Observe that $(1)\rightarrow(2)$ immediately follows from Proposition 3.4. Moreover, for the implication $(2)\rightarrow(1)$, we note that, for any given $d\in\mathbb{Z}_{>0}$, only finitely many components in $\operatorname*{Irr}(A)$ correspond to $A$-modules of dimension $d$. The conclusion follows immediately from Theorem 4.2. ∎
@@ 5. Bricks and generalized standard components
As mentioned previously (see Corollary 3.6 and the paragraph before that), a useful approach to the study of the brick-Brauer-Thrall conjectures is the reduction to quotient algebras. Such reductions are extensively used while showing some of the brick-Brauer-Thrall conjectures for some families (see [CKW] and [MP1, MP2, MP3]). Here we particularly highlight two important points. First, we recall that if $B$ is a quotient algebra of $A$, then $\operatorname*{mod}B$ is a full subcategory of $\operatorname*{mod}A$, thus $\operatorname*{brick}(B,d)\subseteq\operatorname*{brick}(A,d)$, for each $d\in\mathbb{Z}_{>0}$. Second, we note that if $\Gamma_{A}$ has an infinite preprojective component, then there is a quotient algebra of $A$ that is tame concealed (see [SS, Chapter XIV]). Meanwhile, some fundamental results on the (tame) concealed algebras (see [SS, Chapter XII]), together with some geometric facts on bricks and stable modules, imply that all brick-Brauer-Thrall conjectures from Section 2.2 hold for all concealed algebras. Thus, if $\Gamma_{A}$ has a preprojective component, the stable brick conjecture (Conjecture 2.4) is verified for $A$ (compare [CKW, Theorem 3.5 &\&& Conjecture 5.4]), from which the Second brick-Brauer-Thrall conjecture (Conjecture 2.1) will follow for $A$ (compare [Mo1, Section 6]), hence Proposition 3.4 implies that the semibrick conjecture holds for $A$ (Remark 3.5). In contrast, if $\Gamma_{A}$ has an arbitrary (infinite) generalized standard component, $A$ does not necessarily admit a concealed quotient algebra. Thus, we employ a range of geometric and algebraic tools while treating different quotients of algebras under consideration.
@@
Observe that, if $X$ is a homogeneous brick, then $X$ is $\theta_{X}$-stable (see Lemma 2.4).
@@
The above theorem has some interesting applications in the study of the open conjectures from Section 2.2. In particular, the following corollary shows that, for the algebras that admit a generalized standard component, all the conjectures listed in Section 2.2 are equivalent, and more importantly, they all hold for any such algebra.
@@
We finish this section with a short remark. In particular, we recall that $X$ in $\operatorname*{mod}A$ is said to be $\tau$-rigid if $\operatorname{Hom}_{A}(X,\tau X)=0$. Let $\textit{i}\operatorname{\tau-rigid}(A)$ denote the set of all indecomposable $\tau$-rigid indecomposables in $\operatorname*{mod}A$. As shown in [DIJ], for any algebra $A$, the two sets $\textit{i}\operatorname{\tau-rigid}(A)$ and $\operatorname*{brick}(A)$ are closely related, and a good knowledge of either of these sets provides insights to the $\tau$-tilting theory of $A$. In particular, the “brick-$\tau$-rigid correspondence” gives an explicit injective map from $\textit{i}\operatorname{\tau-rigid}(A)$ into $\operatorname*{brick}(A)$, which implies that $A$ is brick-finite if and only if it admits only finitely many indecomposable $\tau$-rigid modules. Consequently, our results on the study of brick and their distribution can be seen through the lense of $\tau$-tilting theory and several other closely related areas of research. Meanwhile, it is known that the maximal number of pairwise Hom-orthogonal components of $\Gamma_{A}$ containing $\tau$-rigid modules is at most $n$ ([Sk1, Corollary 2.2]). Observe that the verbatim brick-analogue of this statement is evidently wrong (consider the Kronecker algebra). However, if $A$ is brick-finite, Proposition 3.1 implies that $\Gamma_{A}$ contains at most $n$ pairwise orthogonal components. We note that the converse of the latter statement is not true in general. More precisely, if $A$ is a wild hereditary algebra, Baer [Ba, Prop. 3.1] showed that if $X$ and $Y$ are two regular modules in $\Gamma_{A}$, then there exists an integer $r$ for which $\operatorname{Hom}_{A}(X,\tau^{r}Y)\neq 0$. Hence, no two regular components in $\Gamma_{A}$ can be orthogonal. Consequently, for a (connected) wild hereditary algebra $A$, although $A$ is evidently brick-infinite, no pair of connected components of $\Gamma_{A}$ are orthogonal.
Acknowledgements. The first-named author was supported by KAKENHI Early-Career Scientist Grant. The second-named author was supported by the Natural Sciences and Engineering Research Council of Canada and by the Canadian Defence Academy Research Programme. The authors would like to thank Lidia Angeleri Hügel, Sota Asai, Osamu Iyama, Rosanna Laking, and Francesco Sentieri for some helpful discussions about problems related to this project.

### v2

@@ 4. Orthogonal orbits and irreducible components
In the following, $A$ is always assumed to be a basic, connected finite dimensional algebra of rank $n$ over an algebraically closed field $k$. For an irreducible component $\mathcal{Z}\in\operatorname*{Irr}(A)$, two orbits $O_{X}$ and $O_{Y}$ in $\mathcal{Z}$ are said to be orthogonal if $X$ and $Y$ are Hom-orthogonal modules. In this section we focus on the behaviour of the Hom-orthogonal modules in the irreducible components of $A$. Through this geometric approach, we obtain a surprisingly elementary equivalent condition for the existence of an infinite family of bricks of the same dimension (Theorem 4.2 and Corollary 4.3). In particular, we prove Proposition B from Section 1. We also give a characterization of brick-finite algebras as a generalization of the local algebras.
We begin with a handy lemma, which should be known to experts, but we have not found it explicitly written in the literature. The proof uses some standard geometric arguments with upper-semicontinuity.
@@
Let $Z$ be such that $\overline{O_{Z}}=\mathcal{Z}$. Assume that the conclusion is false. Then $\mathcal{Z}\times\mathcal{Z}$ contains a pair $(X,Y)$ with $\operatorname{Hom}_{A}(X,Y)=0$. By upper semicontinuity, there is an open (non-empty) set $\mathcal{O}$ in $\mathcal{Z}\times\mathcal{Z}$ for which $(M,N)\in\mathcal{O}$ implies $\operatorname{Hom}_{A}(M,N)=0$. But $O_{Z}\times O_{Z}$ is also an open set of $\mathcal{Z}\times\mathcal{Z}$. Since the latter is irreducible, the open sets $O_{Z}\times O_{Z}$ and $\mathcal{O}$ intersect, which gives the absurdity $\operatorname{Hom}_{A}(Z,Z)=0$. ∎
The preceding lemma leads to the following result, using ideas similar to the proof of Proposition 3.4. In particular, we verify the Strong Semibrick Conjecture for the algebras under consideration (see Remark 3.5).
Proposition 4.2
@@
The previous proposition has some interesting implication. In the following corollary, we give a new characterization of the brick-finite algebras as a generalization of the local algebras from the viewpoint of geometry. Since the result follows from the above theorem, we eliminate the proof. However, to motivate the assertion, let us recall that if $A$ is a local algebra, then for every pair of nonzero $A$-modules $X$ and $Y$, we have $\operatorname{Hom}_{A}(X,Y)\neq 0$.
@@
As another consequence of Proposition 4.2, and in light of Proposition 3.4, we obtain a new equivalent condition for the $2nd$ bBT conjecture. To put this result in perspective, recall that the celebrated Second Brauer-Thrall conjecture (now theorem) asserts that $A$ is a representation-infinite algebra if and only if there are infinitely many $d\in\mathbb{Z}_{>0}$ for which $\operatorname*{ind}(A,d)$ is an infinite set. We observe that, to verify the $2nd$ bBT conjecture for $A$, it suffices to show that for some $d\in\mathbb{Z}_{>0}$ and infinitely many pairs $X$ and $Y$ in $\operatorname*{ind}(A,d)$, we have $\operatorname{Hom}_{A}(X,Y)=0$. In other words, the $2nd$ bBT conjecture asserts $A$ is brick-infinite if and only if there are infinitely many Hom-orthogonal $A$-modules of the same dimension.
@@ 5. Bricks and generalized standard components
As mentioned previously (see Corollary 3.6 and the paragraph before that), a useful approach to the study of the brick-Brauer-Thrall conjectures is the reduction to quotient algebras. Such reductions are extensively used while showing some of the brick-Brauer-Thrall conjectures for some families (see [CKW] and [MP1, MP2, MP3]). Here we particularly highlight two important points. First, we recall that if $B$ is a quotient algebra of $A$, then $\operatorname*{mod}B$ is a full subcategory of $\operatorname*{mod}A$, thus $\operatorname*{brick}(B,d^{\prime})\subseteq\operatorname*{brick}(A,d)$, for each $d\in\mathbb{Z}_{>0}$. Second, we note that if $\Gamma_{A}$ has an infinite preprojective component, then there is a quotient algebra of $A$ that is tame concealed (see [SS, Chapter XIV]). Meanwhile, some fundamental results on the (tame) concealed algebras (see [SS, Chapter XII]), together with some geometric facts on bricks and stable modules, imply that all brick-Brauer-Thrall conjectures from Section 2.2 hold for concealed algebras. Thus, if $\Gamma_{A}$ has a preprojective component, the stable brick conjecture (Conjecture 2.4) is verified for $A$ (compare [CKW, Theorem 3.5 &\&& Conjecture 5.4]), from which the Second brick-Brauer-Thrall conjecture (Conjecture 2.1) will follow for $A$ (compare [Mo1, Section 6]), hence Proposition 3.4 implies that the semibrick conjecture holds for $A$ (Remark 3.5). In contrast, if $\Gamma_{A}$ has an arbitrary (infinite) generalized standard component, $A$ does not necessarily admit a concealed quotient algebra. Thus, we employ a range of geometric and algebraic tools while treating different quotients of algebras under consideration.
@@
Observe that, if $X$ is a homogeneous brick, then $X$ is $\theta_{X}$-stable (see Lemma 2.3). Hence, for homogeneous generalized standard stable tubes, the assertion of the following proposition follows from some earlier observations. However, a direct generalization of the above argument does not seem to work for arbitrary generalized standard stable tubes. Consequently, we construct a stability parameter using the direct sum of some modules along a co-ray ending at the mouth of the tube. This is the core of the following proof.
@@
The above theorem has some interesting applications in the study of the open conjectures from Section 2.2. In particular, the following corollary shows that, for the algebras that admit a generalized standard component, all of the conjectures listed in Section 2.2 are equivalent, and more importantly, they all hold for any such algebra.
@@
Let us finish this section with a remark to highlight some connections between the problems and results in this work and some closely related topics, while pointing out some contrasts. Recall that $X$ in $\operatorname*{mod}A$ is said to be $\tau$-rigid if $\operatorname{Hom}_{A}(X,\tau X)=0$, and let $\textit{i}\operatorname{\tau-rigid}(A)$ denote the set of all indecomposable $\tau$-rigid modules in $\operatorname*{mod}A$. Thanks to some fundamental results from [DIJ], for any algebra $A$, it is known that $\textit{i}\operatorname{\tau-rigid}(A)$ and $\operatorname*{brick}(A)$ are closely related, and a good knowledge of either of these two sets provides insights to the $\tau$-tilting theory of $A$. More specifically, the authors gave a “brick-$\tau$-rigid correspondence”, as an explicit injective map from $\textit{i}\operatorname{\tau-rigid}(A)$ into $\operatorname*{brick}(A)$, and showed that $A$ is brick-finite if and only if $\textit{i}\operatorname{\tau-rigid}(A)$ is a finite set, which is known to be equivalent to the $\tau$-tilting finiteness of $A$. Consequently, our results on the study of bricks and their distribution can be seen through the lense of $\tau$-tilting theory and several other closely related areas of research. Meanwhile, it is known that the maximal number of pairwise Hom-orthogonal components of $\Gamma_{A}$ containing $\tau$-rigid modules is at most $n$ ([Sk1, Corollary 2.2]). Observe that the verbatim brick-analogue of this statement is evidently wrong (consider the Kronecker algebra). However, if $A$ is brick-finite, Proposition 3.1 implies that $\Gamma_{A}$ contains at most $n$ pairwise orthogonal components. Let us remark that the converse of the latter statement is not true in general. More precisely, if $A$ is a wild hereditary algebra, Baer [Ba, Prop. 3.1] showed that if $X$ and $Y$ are two regular modules in $\Gamma_{A}$, then there exists an integer $r$ for which $\operatorname{Hom}_{A}(X,\tau^{r}Y)\neq 0$. Hence, no two regular components in $\Gamma_{A}$ can be orthogonal. Consequently, for a (connected) wild hereditary algebra $A$, although $A$ is evidently brick-infinite, no pair of connected components of $\Gamma_{A}$ are orthogonal.
Acknowledgements. The first-named author was supported by Early-Career Scientist JSPS Kakenhi grant number 24K16908. The second-named author was supported by the Natural Sciences and Engineering Research Council of Canada and by the Canadian Defence Academy Research Programme. The authors would like to thank Lidia Angeleri Hügel, Sota Asai, Osamu Iyama, Rosanna Laking, Francesco Sentieri and Jan Schröer for some helpful discussions about problems related to this project.